Dou is working on a math problem in class, By mistake he multiplies by 10 and divides by 5 instead of multiplying by 5 and then dividing by 10. If the incorrect answer is 3, what is the correct answer?
Answer by Edwin McCravy(20054)

The answer may not be an integer.  Never just assume that every answer has to
come out to be a positive whole number.  You should not rule out negative or
fraction answers unless you can tell that the answer must be a positive whole number.)

>>...By mistake he multiplies by 10 and divides by 5...<<
That's the same as doubling, or multiplying by 2, because 10/5 = 2

>>...instead of multiplying by 5 and then dividing by 10...<<
That's the same as multiplying by 1/2, because 5/10 = 1/2

So he made the mistake of doubling it instead of halving it.  When
he doubled it he got 3.   So he had to have started with half of 3
which is 1 1/2 or 3/2.  So he should have halved that 3/2 and gotten 3/4.

Edwin